OpenStudy (anonymous):

on a vacation, a family drove 1,750 miles over a 10 day period. what was the average miles driven each day?

OpenStudy (dangerousjesse):

What's 50/10?

OpenStudy (anonymous):

5

geerky42 (geerky42):

\[\dfrac{1750~miles}{10 ~days} = \dfrac{?~miles}{1~day}\] So you just simplify 1750/10

OpenStudy (dangerousjesse):

Mhm. now what's 700/10?

geerky42 (geerky42):

you want to have denominator to be 1

OpenStudy (anonymous):

70

OpenStudy (dangerousjesse):

Yep, and 1,000/10?

OpenStudy (anonymous):

100

OpenStudy (dangerousjesse):

Yes ^-^ now what's 5+70+100?

OpenStudy (anonymous):

175

geerky42 (geerky42):

faster way would be "cancel zero" like this: \(\dfrac{1750}{10}\Rightarrow\dfrac{175\cancel{0}}{1\cancel{0}} = 175\)

OpenStudy (anonymous):

thxs jesse u deserve a medal

OpenStudy (dangerousjesse):

So you're answer will be \[\frac{ 175 }{ 1 }\]

OpenStudy (dangerousjesse):

No problem.

OpenStudy (dangerousjesse):

your*

OpenStudy (anonymous):

im on some thing called a plus

OpenStudy (anonymous):

man another hard question

OpenStudy (dangerousjesse):

Just let me know if you need help haha.

OpenStudy (anonymous):

im about to post another question